OpenStudy (anonymous):

Simplify. [2.2 ÷ (4 – 2)] • (–3)

OpenStudy (anonymous):

-3.3 is your answer would u like an explanation as well?

OpenStudy (anonymous):

Yes Proof please.

OpenStudy (anonymous):

Can u help me with another question?

OpenStudy (anonymous):

i can yes

OpenStudy (anonymous):

Compare. Use < or >. –1.1 –0.2

OpenStudy (anonymous):

-1.1 < -0.2

OpenStudy (anonymous):

Thanks :) I have another one if you dont mind..

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

What is the missing number? 2.1 + (–3.4 + 5) = [ + (–3.4)] + 5

OpenStudy (anonymous):

There is a box on the left of the plus sign

OpenStudy (anonymous):

That we have to find the missing number.

OpenStudy (anonymous):

the missing number is 2.1 :)

OpenStudy (anonymous):

Ok last one..

OpenStudy (anonymous):

Which property could be used to find the missing number? (–7.3 + 5.2) + 11 = –7.3 + ( + 11) A. distributive property B. commutative property C. associative property

OpenStudy (anonymous):

it is C. :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

np you want an explanation?

OpenStudy (anonymous):

Yes please.

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

Can u also help me with other questions?

OpenStudy (anonymous):

Associative Property = (a + b) + c = a + (b + c)

OpenStudy (anonymous):

Thats the formula right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok let me give you another questions.

OpenStudy (anonymous):

Describe the meaning of the rate. 6 days each month A. Something happens 6 days per month. B. Each month is 6 days long. C. Each day is one month long. D. A day happens six times each month.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

A. :)

OpenStudy (anonymous):

An airplane travels at an average speed of 250 mph for 3 hours. How far does the airplane travel? A. 83.3 mi B. 250 mi C. 500 mi D. 750 mi

OpenStudy (anonymous):

750 miles D because if it travels at 250 miles per hour then you would multiply 250 by 3 and you get 750

OpenStudy (anonymous):

Wow your really good at this.. Im not best at Mathematics..

OpenStudy (anonymous):

The Peytons' car averaged 66 miles per hour on their trip. The trip took 90 minutes. How far did they travel? A. 90 mi B. 99 mi C. 156 mi D. 5940 mi

OpenStudy (anonymous):

I'm a junior in high school :)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Are you homeschooled or you walk to school?

OpenStudy (anonymous):

90 minutes = 1 hour and 30 mins

OpenStudy (anonymous):

I take virtual school

OpenStudy (anonymous):

you want me to continue to explain or are you good?

OpenStudy (anonymous):

Same aswell, Im in junior high and really am bad at Mathematics..

OpenStudy (anonymous):

Malorie can work 10 math problems in 3 minutes. At this rate, how many problems can she work in one hour? A. 30 B. 60 C. 120 D. 200

OpenStudy (anonymous):

D. 200

OpenStudy (anonymous):

because 60/3 = 20 and then you would multiply 10 and 20 and you get 200

OpenStudy (anonymous):

I wish I was as smart as you :(

OpenStudy (anonymous):

A grocery store is selling 3 granola bars for $1.50. At this rate, how much will 12 granola bars cost? A. $3.00 B. $4.50 C. $6.00 D. $18.00

OpenStudy (anonymous):

Lol just keep practicing and you will get there :P

OpenStudy (anonymous):

Ok xP

OpenStudy (anonymous):

C. $6

OpenStudy (anonymous):

12/3 equals 4 and so you multiply 4 by $1.50 which = 6

OpenStudy (anonymous):

Any more questions my math genius lol?

OpenStudy (anonymous):

ok xd

OpenStudy (anonymous):

so yes?

OpenStudy (anonymous):

Yes 1 second

OpenStudy (anonymous):

LAG XD

OpenStudy (anonymous):

1 sec im getting the question

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

Divide. 1.28 ÷ 0.4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

3.2 :)

OpenStudy (anonymous):

What is the solution to y = 0.7x for the replacement set x = {2}? ( , )

OpenStudy (anonymous):

(2 , 1.4) :)

OpenStudy (anonymous):

So what are you gonna do for summer?

OpenStudy (anonymous):

What is the solution to y = 2x + 3 for the replacement set x = {1.3}? ( , )

OpenStudy (anonymous):

im not sure what the answer is to that i do know that y=5.6

OpenStudy (anonymous):

and im going on a cruise through the Caribbean

OpenStudy (anonymous):

and sorry i have to go

OpenStudy (anonymous):

Bye ;c